OpenStudy (anonymous):

the exact value of cos 2π/8 is? a) radical 3/2 b) - radical 3 / 2 c) 1/2 d) -1/2

OpenStudy (anonymous):

1/rad2

OpenStudy (anonymous):

ya none of those answers are right

OpenStudy (anonymous):

which isn't a option, umm stange. i think your question is wrong

OpenStudy (anonymous):

rationalized denominator is sqrt(2)/2

OpenStudy (anonymous):

but those are the answer that the textbook gives me

OpenStudy (anonymous):

for one thing, \[\frac{2\pi}{8}=\frac{\pi}{4}\]

OpenStudy (anonymous):

yea or cos(45) in degrees

OpenStudy (anonymous):

so maybe there is a typo in the question because as it is written the answer is \[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

wait guys, it's suppose to be 2pi/3

OpenStudy (anonymous):

maybe you didn't type the question correctly, probably the options

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

satellite, come back to my problem

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

wait satellite how do i show work?

OpenStudy (anonymous):

look at a cosine graph

OpenStudy (anonymous):

dockworker how did you get -1/2?

OpenStudy (anonymous):

i mean dockworker is right

OpenStudy (anonymous):

find \[\frac{2\pi}{3}\] on the attached cheat sheet. first coordinate on the unit circle will be cosine it will be \[-\frac{1}{2}\]

OpenStudy (anonymous):

thanks guys :)